A triangle has an angle that measures 10°. The other two angles are in a ratio of 3:14. What are the measures of those two angles?
Answer by ikleyn(52802)

.
The two other angles are 3x and 14x, where x is thoir common measure (now unknown).


Using the fact that the sum of interior angles of any triangles is 180 degrees, you have an equation


    3x + 14x + 10 = 180,    or

    17x = 180 - 10

    17x = 170   =======>  x = 170/17 = 10 degrees.


Then the two unknown angles are 3*10 = 30 degrees  and  14*10 = 140 degrees.


ANSWER.  Two unknown angles are 30 degrees  and  140 degrees.

Solved.